# Principles of Visual Tokens for Efficient Video Understanding

### v1

@@ 5.1 Comparison to State-of-the-art
Tables 1 (also shown in Fig. 1) and 2 show the results of LITE on SS-V2 and Kinetics-400 respectively. We make several observations.
First, LITE offers a better trade-off between computational cost and accuracy than prior methods. At fixed accuracy, LITE requires less compute, and at fixed compute, it achieves higher accuracy. For example, on the SS-V2 dataset, LITE uses only 40% of ObjectViViT [47]’s GFLOPs with similar accuracy. On Kinetics-400, LITE achieves comparable accuracy with 12% of the GFLOPs used by LookupViViT [20]. LITE also outperforms token-merging methods: compared to ToMe [4], it achieves similar GFLOPs reduction with an accuracy gain of  0.3% on SS-V2 and 0.5% on Kinetics-400. Likewise, LITE slightly outperforms STA [9] at equivalent GFLOPs.
@@ 5.3 Analysis of LITE
Model Top-1 Top-5
Simple MLP selector 65.03 88.52
Edges 64.98 88.92
Isolated peaks 64.83 88.67
Sharpen distribution 64.80 87.98
Global branch 65.10 88.55
Table 5: Results of impact of different variants of data cleaning strategies and integration of global branch. Results are tested with 4K samples of the SS-V2 test set.
Oracle creation. Gradient scores in the oracle can be calculated in two different ways: either using the true label or using the highest predicted class. Table 3 displays the performance of these two variants keeping 50% of tokens. We observe that using the true label provides a very large improvement of 8.5%.
@@
Variants. It is remarkable to see in Table 5 that none of the reasonable ideas from Sec. 4.4 significantly outperform a simple MLP. This points to a striking conclusion: the oracle is extremely hard to predict, and the MLP achieves a balance between accuracy and avoiding overfitting.
@@
TimeSformer-LITE 0.1 15 11.68 29.54
Table 6: Results of LITE on a different backbone on SS-V2 dataset. While LITE is still able to retain accuracy in different backbones, the quality of the oracle is significantly worse, affecting the overall performance.

### v2

@@ 5.1 Comparison to State-of-the-art
Tables 1 (also shown in Fig. 1) and 2 show the results of the balance between GFLOPs and Top-1 accuracy for LITE and state-of-the-art on SS-V2 and Kinetics-400 respectively. We make several observations.
First, LITE offers a better trade-off between computational cost and accuracy than prior methods. In other words: for a fixed accuracy, LITE requires less compute, and for a fixed compute, it achieves higher accuracy. For example, on the SS-V2 dataset, LITE uses only 40% of ObjectViViT [53]’s GFLOPs with similar accuracy. On Kinetics-400, LITE achieves comparable accuracy with 12% of the GFLOPs used by LookupViViT [23]. LITE also outperforms token-merging methods: compared to ToMe [4], it achieves similar GFLOPs reduction with an accuracy gain of  0.3% on SS-V2 and 0.5% on Kinetics-400. Likewise, LITE slightly outperforms STA [11] at equivalent GFLOPs.
@@
5.2 Zero-shot across Datasets and Tasks
To evaluate the generalization capability of LITE , we conduct two sets of experiments. First, we test LITE trained on Kinetics-400 on SS-V2 and UCF101 to assess its ability to generalize across different datasets within the same task. The results of Table 3 demonstrate LITE’s strong zero-shot capability, even to SS-V2 which is visually very different from Kinetics-400. Compared to our results in Table 1, we observe that LITE exhibits strong cross-domain generalization, achieving performance close to that of a selector trained from scratch. Second, we examine the performance of the same LITE model trained on classification on an even more challenging setting: an unseen task (action detection) and unseen dataset (AVA) to evaluate its cross-task transferability. Table 3 highlights LITE’s strong cross-task generalization ability, reducing GFLOPs by 35% while maintaining a minimal 0.7% drop in mAP.
Test Train Model GFLOPs Top-1 / mAP
Classification UCF101 Classification K400 VideoMAE [39] 181 91.25
[…]
Classification SS-V2 Classification K400 VideoMAE [39] 181 70.50
[…]
Detection AVA Classification K400 VideoMAE [39] 181 31.50
[…]
Table 3: Generalization performance of the selector trained on Kinetics-400 and tested on the different datasets and tasks. The results of VideoMAE serve as baseline, which are trained and tested on the same dataset. We use Top-1 accuracy as the metric for classification task and mAP for detection task.
@@ 5.4 Analysis of LITE
Oracle creation. Gradient scores in the oracle can be calculated in two ways: using the true label or the highest predicted class. Table 4 shows their performance with 50% of tokens. We observe that using the true label provides a very large improvement of 8.5%.
@@
Table 6: LITE’s results on a different backbone on SS-V2 dataset.